OpenStudy (anonymous):

5v^2+31v-28 how do you factor?

OpenStudy (antonio_xx2):

7

OpenStudy (anonymous):

@antonio_xx2 could you show me how to do it?

OpenStudy (anonymous):

You need to find two numbers that when multiplied give 28 and when added give 31.

OpenStudy (anonymous):

what about the five?

OpenStudy (anonymous):

what numbers even factor out like tht?

OpenStudy (anonymous):

With numbers like these I would suggest to just solve it with the quadratic fomula.

OpenStudy (anonymous):

right, thts what im doing right now

OpenStudy (anonymous):

but which numbers do tht?

OpenStudy (anonymous):

See what 2 results it gives you, and then put their opposites in this form: \[(v \pm x)(v \pm x)\] (where x are the results you find)

OpenStudy (anonymous):

what is tht?

OpenStudy (anonymous):

What did the quadratic formula yield?

OpenStudy (anonymous):

i dont know how to do tht

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2-4 \times a \times c} }{ 2 \times a }\]

OpenStudy (anonymous):

In your case a = 5, b = 31, c = -28

OpenStudy (anonymous):

im not doing it tht way for school, i dont even know where you're goin with tht

OpenStudy (anonymous):

Did you do similar exercises?

OpenStudy (anonymous):

On other equations?

OpenStudy (anonymous):

well for r^2-r-6 i got r=2 r=-3

OpenStudy (anonymous):

Do this: x^2+15x+50

OpenStudy (anonymous):

x=10 x=5

OpenStudy (anonymous):

It's supposed to be easy on equations like this, but the one you gave at the beginning is only possible by someone with a relatively high maths skill.

OpenStudy (anonymous):

well my math teacher gave them too me and taught us how to do them, but its some asktik thing but it's really confusing

OpenStudy (anonymous):

Which is why I suggested using the quadratic formula, it's sort of the last resort if you can't find two numbers.

OpenStudy (anonymous):

Go above, and plug in the numbers for me, tell me what you get.

OpenStudy (anonymous):

oh ok then, let me copy tht down on my paper

OpenStudy (anonymous):

what does tht funky plus sign mean?

OpenStudy (anonymous):

It means do one time plus, and on another a minus.

OpenStudy (anonymous):

Which also means you will get two answers.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

did you reach -31plusminus39 / 10 ?

OpenStudy (anonymous):

i dont know how to use tht sign

OpenStudy (anonymous):